Please solve this formula for the given variable.
b+3/e=f/2
solve for e
Please show your work.
Thanks!

Answer by checkley77(12844)

b+3/e=f/2
(be+3)/e=f/2
now cross multiply.
ef=2(be+3)
ef=2be+6
ef-2be=6
e(f-2b)=6
e=6/(f-2b) ans.